# DehazeGS: Seeing Through Fog with 3D Gaussian Splatting

### v5

@@ Introduction
Baselines. In addition to comparing with the original 3DGS (Kerbl et al. 2023), we also compared our method with ScatterNeRF (Ramazzina et al. 2023) and DehazeNeRF (Chen et al. 2023). We further evaluated the reconstruction performance of combining low-level computer vision dehazing models with 3DGS. Specifically, we first dehazed the multi-view images using the pretrained models of Dehamer (Guo et al. 2022) or DehazeFormer (Song et al. 2023) (as shown in the supplementary materials), which were then fed into 3DGS for reconstruction. Furthermore, since SeaSplat (Yang, Leonard, and Girdhar 2024) (a real-time rendering method for underwater scenes) also has some dehazing reconstruction capabilities, we included it in the comparison as well. Since DehazeNeRF has not released its source code, its qualitative and quantitative results are both sourced from its original paper. Note that the views of the real dataset presented in DehazeNeRF were specifically cropped (removing some edge regions of the scene). To ensure a fair comparison, we also cropped the edge regions of the rendered views accordingly.
Training Details. ScatterNeRF was trained on three 4090 GPUs with a batch size of 2048. All other models were trained on a single 4090 GPU, except for DehazeNeRF, whose results are taken from its original paper. SeaSplat was trained on a single GPU. All other parameters followed the default settings. For our model, training was performed on a single GPU. The number of iterations was set to 30K for synthetic datasets and 3K for real-world datasets. $\lambda_{D}$, $\lambda_{B}$, and $\lambda_{d_{rec}}$ were all set to 0.1. The learning rate for the convolutional network was set to $1\times 10^{-7}$, with a weight decay factor of 0.1, decaying over the corresponding number of iterations for synthetic and real scenes.

### v6

@@ Introduction
Baselines. In addition to comparing with the original 3DGS (Kerbl et al. 2023), we also compared our method with ScatterNeRF (Ramazzina et al. 2023) and DehazeNeRF (Chen et al. 2023). We further evaluated the reconstruction performance of combining low-level computer vision dehazing models with 3DGS. Specifically, we first dehazed the multi-view images using the pretrained models of Dehamer (Guo et al. 2022) or DehazeFormer (Song et al. 2023) (as shown in the supplementary materials), which were then fed into 3DGS for reconstruction. Furthermore, since SeaSplat (Yang, Leonard, and Girdhar 2024) (a real-time underwater rendering method) also has some dehazing reconstruction capabilities, we included it in the comparison as well. Since DehazeNeRF has not released its source code, its qualitative and quantitative results are both sourced from its original paper—for the Lion and Bear scenes (where DehazeNeRF lacks quantitative data), we calculated such results for corresponding viewpoints for fair comparison. Note that the real dataset views in DehazeNeRF were specifically cropped (removing scene edges). For fair comparison, we similarly cropped edges of our rendered views.
Training Details. ScatterNeRF was trained on three 4090 GPUs with a batch size of 2048. All other models were trained on a single 4090 GPU (except for DehazeNeRF). SeaSplat was trained on a single GPU. All other parameters followed the default settings. For our model, training was performed on a single GPU. The number of iterations was set to 30K for synthetic datasets and 3K for real-world datasets. $\lambda_{D}$, $\lambda_{B}$, and $\lambda_{d_{rec}}$ were all set to 0.1. The learning rate for the convolutional network was set to $1\times 10^{-7}$, with a weight decay factor of 0.1, decaying over the corresponding number of iterations for synthetic and real scenes.